### v2

# Macroscopic degeneracy of ground state in the frustrated Heisenberg diamond chain

Abstract
The spin-$\frac{1}{2}$ Heisenberg diamond chain with ferro- and antiferromagnetic exchange interactions is studied. The phase boundary in the parametric space of these interactions is determined, where the transition between the ferromagnetic and other (singlet or ferrimagnetic) ground state phases occurs. On this phase boundary there is a dispersionless (flat) energy band in the one-magnon spectrum and these states can be represented as localized magnons in the trapping cells between neighboring diamonds. The ground state consists of the localized magnons and special magnon complexes and it is macroscopically degenerate. A remarkable feature of the model is the existence, on a certain part of the phase boundary, of two- and three-magnon localized states, forming flat bands in two- and three-magnon spectrum. All these localized states also belong to the ground state manifold, which turns out to be exactly the same as for the system of independent spins-$\frac{3}{2}$. This implies a large macroscopic degeneracy of the ground state $4^{n}$ ($n$ is number of diamonds in the chain) and a high residual entropy per spin $s_{0}=\frac{2}{3}\ln 2$.
I Introduction
The low-dimensional quantum magnets on geometrically frustrated lattices have been extensively studied during the last years [1, 2]. There is a peculiar class of frustrated quantum magnets in which the ground state is macroscopically degenerate. Systems with macroscopic degeneracy have attracted great interest and are usually described by models with flat bands [3, 4]. Flat-band in the one-magnon spectrum of the frustrated Heisenberg spin system means that the magnons are localized in a small portion of the lattice (trapping cells). The localization of the magnons arises due to a destructive quantum interference, caused by frustration. There are highly frustrated antiferromagnetic spin systems which support a completely dispersionless magnon band [4, 5, 6, *zhit2, 8], so that the excitations in this band are localized states. The localization of one-magnon states is a base for the construction of multi-magnon states, because a state consisting of independent (non-overlapping) localized magnons is an exact eigenstate. Such systems include, for example, the delta-chain, the kagome lattice, kagome-like chains, the Tasaki lattice etc. Their important feature is the triangular geometry of antiferromagnetic bonds. The ground state of such systems at the saturation magnetic field consists of independent localized magnons and it is macroscopically degenerate [4]. The ground state and low-temperature properties for the antiferromagnetic Heisenberg models with flat band have been actively studied over the last decades. It was shown that flat band physics may lead to new interesting phenomena such as the residual entropy at the saturation magnetic field, the zero-temperature magnetization plateau and the magnetization jump, an extra low-temperature peak in the specific heat etc. [9, 10, 11, 8, 12]. Such systems have enhanced magnetocaloric effect and are promising for efficient low-temperature cooling [10, 11, 13, 12].
Figure 1: Delta chain spin model. Shaded area indicates the trapping valley.
Recently it was found that the localized magnon states can exist in a certain frustrated spin system with competing ferro- (F) and antiferromagnetic (AF) interactions. An example of such model is spin-$\frac{1}{2}$ Heisenberg delta-chain with the ferromagnetic $J_{1}<0$ and the antiferromagnetic $J_{2}>0$ interactions (see Fig.1). For $J_{2}=-\frac{J_{1}}{2}$ the localized states exist in this model. However, in contrast to the AF-AF delta-chain, the localized magnon states of the F-AF chain are exact ground states at zero magnetic field. Besides, the ground state manifold contains the special states with overlapping magnons (localized multi-magnon complexes). Thus, the ground state degeneracy in this model is even higher than the one for the AF-AF delta chain. The properties of such F-AF delta chain have been studied in Refs.[14, 15, *DKRS2, *DKRS3].
Figure 2: Diamond chain spin model. Shaded area indicates the trapping cell.
In this paper we consider one more example of the model, in which the localized magnons lead to the macroscopic degeneracy of the ground state. This is the spin-$\frac{1}{2}$ Heisenberg diamond chain, shown in Fig.2. Generally, this model is one of the examples where the interplay of quantum effects and frustration leads to a wide variety of ground state phases; it has attracted a lot of attention in both experimental and theoretical studies. It was proposed as a minimal theoretical model of several copper-based compounds as azurite $Cu_{3}(CO_{3})_{2}(OH)_{2}$ [18] and alumoklyuchevskite $K_{3}Cu_{3}AlO_{2}(SO_{4})_{4}$ [19]. Besides, the diamond model is also interesting in its own right. The ground state phase diagram of the diamond chain has been extensively studied in many works [20, 21, 22, 23, 24, 25, 26, 27, 28, 29] and it was shown that the diamond model exhibits various quantum phases depending on exchange interactions. In particular, the $1/3$ magnetization plateau was determined for certain ratios between exchange interactions [30, 31, 28, 27, 29].
In this work we will study the spin-$\frac{1}{2}$ Heisenberg diamond chain with F and AF competing interactions. We will show that under special relations between exchange interactions, describing the boundary between the ferromagnetic and other (singlet or ferrimagnetic) phases, the ground state is macroscopically degenerate.
The paper is organized as follows. In Sec.II the model Hamiltonian with generally different exchange interactions is presented and the conditions for the interactions in which the ground state has macroscopic degeneracy are determined. It will be shown that, under certain conditions, along with the localized one-magnon states there are also localized two- and three-magnon states. We present four types of the diamond chain with different ground state degeneracy. In the first type only one-magnon localized states exist. In the second one there are one-, two- and three- magnons in the trapping cell and this case includes the diamond chain with equivalent opposite sides of a diamond (we name it as the ‘distorted’ diamond chain). The third model is the ideal (symmetric) diamond chain and in the fourth type of the model one-magnon states are localized on diagonals of diamonds. The ground state degeneracy for these four types of diamond chain models are studied in Subsecs. IIA, IIB, IIC and IID. In concluding Section we give a summary of our results.
II Diamond chain at the critical point
The Hamiltonian of the spin $s=\frac{1}{2}$ Heisenberg diamond chain can be represented as a sum of local Hamiltonians
$\hat{H}=\sum_{i=1}^{n}\hat{H}_{i}$ (1)
and $\hat{H}_{i}$ is the Hamiltonian of $i$-th diamond
$\hat{H}_{i}=J_{1}\mathbf{s}_{i}\cdot\mathbf{\xi}_{i}+J_{2}\mathbf{\xi}_{i}\cdot\mathbf{\eta}_{i}+J_{3}\mathbf{s}_{i}\cdot\mathbf{\eta}_{i}+J_{4}\mathbf{s}_{i+1}\cdot\mathbf{\xi}_{i}+J_{5}\mathbf{s}_{i+1}\cdot\mathbf{\eta}_{i}-\frac{J_{0}}{4}$ (2)
where $\mathbf{s}_{i}$, $\mathbf{\xi}_{i}$, $\mathbf{\eta}_{i}$ are spin operators of $i$-th diamond, $n$ is total number of diamonds and $J_{0}=J_{1}+J_{2}+J_{3}+J_{4}+J_{5}$. In general, there are five different exchange interactions $J_{i}$ as shown in Fig.2. The constant in Eq.(2) is chosen so that the energy of the state with the maximal spin of diamond ($S=2$) is zero.
It is known [14] that the macroscopic ground state degeneracy in frustrated spin systems with competing (F) and (AF) interactions arises if the local Hamiltonian has several degenerate ground states, one of which is the state with the maximal spin. Four spins of diamond form one quintet ($S=2$), three triplets ($S=1$) and two singlets ($S=0$). The condition that the quintet and one of three triplets are degenerate with zero energy is
$J_{2}=-\frac{J_{1}J_{3}}{J_{1}+J_{3}}-\frac{J_{4}J_{5}}{J_{5}+J_{4}}$ (3)
The conditions that these eight states (5 states of quintet and 3 states of triplet) are the ground states with $E_{i}=0$ of $\hat{H}_{i}$ and all other eigenvalues $E_{i}>0$ are
$J_{1}+J_{3}<0,\qquad J_{4}+J_{5}<0$ (4)
Relations (3) and (4) determine the range of interaction values corresponding to the transition point between the ferromagnetic and other (singlet or ferrimagnetic) ground state phases of model (1). For simplicity we chose the interaction $J_{2}$ as a parameter defining the transition point. Eq.(3) gives the critical value of this parameter. For $J_{2}$ less than the value given by Eq.(3), the ground state of model (1) is ferromagnetic. Further we assume that both relations (3) and (4) are satisfied.
Because the neighboring local Hamiltonians $\hat{H}_{i}$ do not commute with each other, the following inequality for the lowest eigenvalue $E_{0}$ of $\hat{H}$ is valid:
$E_{0}\geq\sum E_{i}=0$ (5)
The energy of the ferromagnetic state of $\hat{H}$ with maximal total spin $S_{\max}=\frac{3n}{2}$ is zero. Therefore, the inequality in (5) turns into an equality and the ground state energy is zero.
The relations (3) and (4) can also be obtained alternatively from the condition that a one-magnon excitation band with the spin $S^{z}=S_{\max}^{z}-1$ is dispersionless (flat band) with zero energy. The dispersionless one-magnon states correspond to localized states which can be chosen as
$\hat{\varphi}_{1i}\left|F\right\rangle=(\sigma_{i-1}^{-}+s_{i}^{-}+\tau_{i}^{-})\left|F\right\rangle\text{ \qquad}i=1,2,\ldots n$ (6)
where $\left|F\right\rangle$ is the ferromagnetic state with all spins up, $s_{i}^{-}$ are lowering spin operators and new convenient operators $\sigma_{i}^{-}$, $\tau_{i}^{-}$ relate to lowering spin operators $\eta_{i}^{-},\xi_{i}^{-}$ as follows:
$\sigma_{i}^{-}$ $=$ $\frac{J_{4}+J_{5}}{J_{1}J_{5}-J_{3}J_{4}}(J_{1}\eta_{i}^{-}-J_{3}\xi_{i}^{-})$ (7)
$\tau_{i}^{-}$ $=$ $-\frac{J_{1}+J_{3}}{J_{1}J_{5}-J_{3}J_{4}}(J_{4}\eta_{i}^{-}-J_{5}\xi_{i}^{-})$ (8)
so that $\sigma_{i}^{-}+\tau_{i}^{-}=\eta_{i}^{-}+\xi_{i}^{-}$.
The states $\hat{\varphi}_{1i}\left|F\right\rangle$ are localized in the trapping cells between neighboring diamonds as shown in Fig.2 by a shaded area. All $n$ states (6) are the ground states and they are linearly independent, because each $\hat{\varphi}_{1i}$ includes the corresponding operator $s_{i}^{-}$, which is not shared with other functions $\hat{\varphi}_{1j}$.
There are three one-magnon bands of model (1), one of them is flat, and the energies of the other two bands are given by
$2E(q)=-J_{0}\pm\sqrt{J_{0}^{2}-3J_{2}(J_{0}-J_{2})-3(J_{1}+J_{4})(J_{3}-J_{5})-2(J_{3}J_{5}+J_{1}J_{4})(1-\cos q)}$ (9)
It can be shown that the energy $E(q)>0$ if both conditions (3) and (4) are satisfied.
As follows from Eqs.(7),(8), there is one special case in which the denominator in these equations goes to zero:
$\frac{J_{3}}{J_{1}}=\frac{J_{5}}{J_{4}}$ (10)
In this special case the one-magnon states (6) are localized on diagonals of diamonds and can be written as
$\hat{\phi}_{i}\left|F\right\rangle=(J_{1}\eta_{i}^{-}-J_{3}\xi_{i}^{-})\left|F\right\rangle\text{ \qquad}i=1,2,\ldots n$ (11)
Therefore, we name this special case as ‘diagonal state case’, the properties of the model for this special case will be studied separately in Sec.IId.
A remarkable feature of model (1) is the fact that along with the presence of one-magnon localized states, under certain condition, there are exact two- and three-magnon localized states with zero energy. These localized states form flat-bands in two- and three-magnon spectrum. Omitting all intermediate calculations we give the final expressions for two- and three-magnon states, localized in the $i$-th trapping cell (Fig.2):
$\hat{\varphi}_{2i}\left|F\right\rangle$ $=$ $(\sigma_{i-1}^{-}s_{i}^{-}+s_{i}^{-}\tau_{i}^{-}+\sigma_{i-1}^{-}\tau_{i}^{-})\left|F\right\rangle$ (12)
$\hat{\varphi}_{3i}\left|F\right\rangle$ $=$ $\sigma_{i-1}^{-}s_{i}^{-}\tau_{i}^{-}\left|F\right\rangle$ (13)
It can be checked that Eqs.(12),(13) give exact two- and three-magnon eigenstates of the Hamiltonian (1) with zero energy if the exchange interactions satisfy the relation
$\frac{J_{3}}{J_{1}}=\frac{J_{4}}{J_{5}}$ (14)
Then, two- and three-magnon states (12) and (13) also belong to the ground state manifold. It is worth noting that when the relation (14) is valid, the ground state degeneracy of each local Hamiltonian $\hat{H}_{i}$ increases. In this case, one singlet state is added to 8 degenerate ground states with $S=2$ and $S=1$.
Relation (14) is satisfied automatically for the important case of the ‘distorted’ diamond chain, in which $J_{3}=J_{4}$ and $J_{1}=J_{5}$ (but $J_{1}\neq J_{3}$).
There is one more special case, when both relations (10) and (14) are satisfied simultaneously. In this case $J_{1}=J_{3}=J_{4}=$ $J_{5}=-1$ and $J_{2}=1$ (ideal diamond chain) and localized one-magnon functions are diagonal singlets $(\eta_{i}^{-}-\xi_{i}^{-})\left|F\right\rangle$. As it will be shown below, the ground state degeneracy is maximal in this case.
Figure 3: Examples of four different types of diamond spin chains with macroscopic ground state degeneracy. The values of the exchange integrals are written close to the corresponding bonds. For clarity, the ferromagnetic bonds (all with $J=-1$) are shown by solid lines, antiferromagnetic bonds - by dashed lines.
Thus, we have determined the necessary conditions (3) and (4) for the macroscopic ground state degeneracy of the diamond chain model (1). Based on the above analysis we can distinguish four different types of the diamond chain with the macroscopic ground state degeneracy, depending on whether the relations (10) and (14) are satisfied or not. Examples of four types of diamonds are shown in Fig.3. We will name these types as follows: the ‘general case’ is the one when none of the relations (10) and (14) are fulfilled (Fig.3a); the ‘distorted diamond case’ is realized when the relation (14) is satisfied, but Eq.(10) is not (Fig.3b); the ‘ideal diamond case’ is when both relations are fulfilled (Fig.3c); and the ‘diagonal diamond case’ is when the relation (10) is satisfied, but Eq.(14) is not (Fig.3d).
Figure 4: Phase diagram of spin diamond chain (1) with diagonal bond interaction $J_{2}$ given by Eq.(3). The special cases are shown by solid line (‘distorted diamond case’), dashed line (‘diagonal state case’) and by solid circle (‘ideal diamond case’). Thick open circles denote four particular cases of the diamond model shown in Fig.3 and they are signed with the corresponding letters a, b, c and d.
As it follows from the relations (10) and (14), it is convenient to plot the phase diagram of the model (1) in the $J_{3}/J_{1}$ and $J_{5}/J_{4}$ axes (see Fig.4). In Fig.4 the ‘diagonal state case’ is shown by a dashed line, the distorted diamond case by a solid line and the ideal case by a solid circle. The entire plane, except for these special cases, corresponds to the general case. However, in order to satisfy the conditions (4), there is a prescription on how to choose the signs of interactions $J_{i}$ corresponding to a definite point on the phase diagram Fig.4. If the point is in the region $J_{3}/J_{1}>0$, then the condition $J_{1}+J_{3}<0$ leads uniquely to the conclusion that both $J_{3}$ and $J_{1}$ are ferromagnetic ($J_{3}$, $J_{1}<0$). If $-1<J_{3}/J_{1}<0$, then $J_{3}>0$ and $J_{1}<0$. And, finally, if $J_{3}/J_{1}<-1$, then $J_{3}<0$ and $J_{1}>0$. Similar conditions are applied for the correct choice of signs of interactions $J_{4}$ and $J_{5}$ in different regions of the phase diagram. To illustrate, four particular cases of the diamond model shown in Fig.3, are represented by thick open circles in Fig.4 and signed with the corresponding letters: a, b, c and d.
II.1 General case
Let us first consider the case when both relations (10) and (14) are not satisfied. The example of the diamond of this type is shown in Fig.3a, where the exchange integrals are chosen so that the diagonal interaction $J_{2}$ is zero. In the general case there are only one-magnon localized states with zero energy as in the F-AF delta-chain at the critical value of the frustration parameter. In the latter case the one-magnon state with zero energy is localized in a valley between neighboring triangles (see Fig.1). Any state consisting of independent (isolated) magnons has zero energy and belongs to the ground state manifold. As it is shown in [14] for the F-AF delta-chain, in addition to the independent localized states there are also specifically overlapping magnons in the ground states. For the general case of the diamond chain the problem of counting of these states is completely similar to the one for the F-AF delta-chain, which was solved in [14]. Rewriting these results to the case of the diamond chain with periodic boundary conditions (PBC) we find that the number of ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is
$W(n,k)$ $=$ $C_{n}^{k}\qquad 0\leq k\leq n/2$ (15)
$W(n,k)$ $=$ $C_{n}^{n/2}\qquad n/2+1\leq k\leq 3n/2$ (16)
where $C_{n}^{k}=\frac{n!}{k!(n-k)!}$ are binomial coefficients.
The total number of the ground states $W(n)$ is
$W(n)=2^{n}+(n+4)C_{n}^{n/2}$ (17)
According to this equation the ground state degeneracy is exponentially large. It leads to the finite residual entropy per spin (total number of spins is $N=3n$)
$s_{0}=\frac{\ln W(n)}{N}$ (18)
and $s_{0}=\frac{1}{3}\ln 2$ at $n\gg 1$.
Figure 5: Dependence of the energy gap (logarithmic scale) on the system size for the general case (Fig.3a), calculated numerically for $N=12,18,24$ (exact diagonalization and Lanczos algorithm).
The properties of the excitation spectrum of the diamond chain in the general case are similar to that for the F-AF delta-chain. Numerical calculations show that the gap in the spectrum is exponentially small at $n\gg 1$ as in the F-AF delta-chain (see Fig.5).
Figure 6: Temperature dependencies of the specific heat (solid line), entropy (dashed line) and the product of magnetic susceptibility and temperature, $\chi T$, (dotted line) for the general case (Fig.3a), calculated numerically for $N=18$ (exact diagonalization). For better visibility the specific heat is multiplied by a factor of 3, $3C(T)$.
The low-temperature thermodynamic properties in this case are determined by these low-lying excitations and they lead to the specific behavior of the thermodynamic quantities at low temperature. For example, $C(T)$ has low-temperature maximum and a long low-temperature tail, as shown in Fig.6. The entropy in Fig.6 tends to the finite value in accordance with Eq.(18) at $T\to 0$ and increases to the value $s_{0}=\ln 2$ at $T\to\infty$.
The behavior of the magnetization and the uniform susceptibility of the diamond chain (in the general case) is similar to the behavior of the F-AF delta-chain, and this point was discussed in detail in [14]. We give here a brief summary of the results [14].
Since there are exponentially low excited states, the contribution of excited states to magnetization and susceptibility can not be neglected even at very low temperatures. Nevertheless, it is useful to consider the magnetization given only by the contribution of the ground state manifold. This magnetization $m_{0}$ is a function of variable $x=\frac{h}{T}$. At $x\gg\frac{1}{N}$ the magnetization $m_{0}$ for the general case of diamond chain is
$m=\frac{M}{N}=\frac{1}{2}-\frac{1}{3(1+\exp(h/T))}$ (19)
Figure 7: Magnetization curve as a function of $h/T$ for the general case (Fig.3a), calculated numerically for $N=18$ and $T=0.001$ (thick solid line) and given by Eq.(19) (thin solid line).
The magnetization curve (19) is shown in Fig.7 together with the magnetization curve calculated numerically for $N=18$ and $T=0.001$. As one can see in Fig.7, the calculated curve coincides with the curve (19) for $h/T>1$, but the limit $h/T\to 0$ is different. The reason of this discrepancy is related to the singular behavior of magnetization and susceptibility in low magnetic fields.
The numerically calculated magnetization is zero at $h=0$ and increases linearly with increasing magnetic field, $m=\chi h$. The magnetic susceptibility $\chi$ depends on the temperature and the system size. An analysis similar to the one given in [14] shows that in the thermodynamic limit the magnetic susceptibility diverges at $T\to 0$ as $\chi\sim T^{-\alpha}$ with the exponent $\alpha\approx 1.25$. Thus, the true magnetization curve in the thermodynamic limit, $m(h)$, starts at $m(0)=0$, then increases linearly with a very steep slope,
$m\sim\frac{h}{T^{\alpha}}$ (20)
so that it approaches the value $m=1/3$ at $h_{c}\sim T^{\alpha}$. For $h>h_{c}$ the true magnetization is described by Eq.(19).
When the magnetization curve is plotted as a function of $h/T$, as shown in Fig.7, the low-field region where the true magnetization curve deviates from Eq.(19) is limited by $h_{c}/T\sim T^{\alpha-1}$, so that this region vanishes in the limit $T\to 0$. This implies that the magnetization at $\frac{h}{T}\to 0$ tends to $\frac{1}{3}$. Therefore, the diamond chain in the general case is magnetically ordered at the transition point at $T=0$ and, therefore, the magnetization undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point.
II.2 ‘Distorted’ diamond chain
Let us consider the ‘distorted’ diamond chain in which the trapping cell can contain localized one-, two- and three-magnon states. In this case condition (14) is satisfied but (10) is not. Further in this subsection we focus on the most interesting from the experimental point of view case $J_{3}=J_{4}$ and $J_{1}=J_{5}$ (but assuming $J_{1}\neq J_{3}$). A particular case of such choice corresponds to the distorted diamond chain shown in Fig.3b. In fact, this is a generic case for the the ‘distorted’ diamond chain, so that the results for the case $J_{3}\neq J_{4}$, $J_{1}\neq J_{5}$ are qualitatively similar.
All possible states consisting of independent (isolated) magnons have zero energy and belong to the ground states. However, in similarity to the previous case such states do not exhaust all possible ones with zero energy. For example, definite combinations of overlapping magnons are also the exact ground states. For the F-AF delta chain with open boundary condition (OBC) such multi-magnon states can be chosen as products
$\Phi_{1}(m_{1})\Phi_{1}(m_{2})\ldots\Phi_{1}(m_{k})\left|F\right\rangle\qquad 1\leq m_{1}<m_{2}<\ldots<m_{k}\leq n-1$ (21)
where
$\Phi_{1}(m)=2S^{-}(m)+\eta_{m+1}^{-}$ (22)
and $S^{-}(m)$ is the total lowering spin operator of all spins for the first $m$ triangles and $\eta_{m+1}^{-}$ is the lowering operator of $(m+1)$-th apical spin.
As it was shown in [14], the product (21) is the exact eigenfunction of each triangle Hamiltonian including $m$-th and $(m+1)$-th triangles.
For the diamond chain the analog of the operator $\Phi_{1}(m)$ has the form
$\Phi_{1}(m)=S^{-}(m)+\tau_{m+1}^{-}$ (23)
where $\tau_{m+1}^{-}$ is provided in Eq.(8) and
$S^{-}(m)=s_{m+1}^{-}+\sum_{i=1}^{m}(s_{i}^{-}+\xi_{i}^{-}+\eta_{i}^{-})$ (24)
For the diamond chain with two- and three-magnons localized in the same trapping cell this approach can be extended to include functions similar to $\Phi_{1}(m)$ for two- and three-magnon states. These functions have the forms
$\Phi_{2}(m)$ $=$ $S^{-}(m)[\frac{1}{2}S^{-}(m)+\tau_{m+1}^{-}]$ (25)
$\Phi_{3}(m)$ $=$ $S^{2-}(m)[\frac{1}{3}S^{-}(m)+\tau_{m+1}^{-}]$ (26)
It can be checked that the functions $\Phi_{\sigma}(m)\left|F\right\rangle$ ($\sigma=1,2,3$) are exact eigenfunctions of all local Hamiltonians with zero energy. Then we can construct the following products for the diamond model with (OBC)
$\Phi_{\sigma_{1}}(m_{1})\Phi_{\sigma_{2}}(m_{2})\ldots\Phi_{\sigma_{k}}(m_{k})\left|F\right\rangle\text{ \ \ }1\leq m_{1}<m_{2}<\ldots<m_{k}\leq n-1$ (27)
where $\sigma_{i}=1,2,3$.
These products are the zero-energy eigenfunctions of each local Hamiltonian and they form the ground states. We can interpret these products as the states of the system of $n$ non-interacting spins $\frac{3}{2}$. For example, let us consider the product
$\Phi_{1}(m_{1})\Phi_{3}(m_{2})\Phi_{2}(m_{3})\Phi_{2}(m_{4})\left|F\right\rangle$ (28)
This product can be identified as the state of the system with spin $s=\frac{3}{2}$ consisting of $(n-4)$ spins with $s^{z}=\frac{3}{2}$, one spin with $s^{z}=\frac{1}{2}$, two spins with $s^{z}=-\frac{1}{2}$ and one spin with $s^{z}=-\frac{3}{2}$. Generally, the number of ground states of the diamond model in the spin sector $S^{z}$ coincides with that for the system of non-interacting spins $\frac{3}{2}$.
Further, we consider the case of the periodic boundary conditions (PBC). Any state consisting of isolated magnons
$\varphi_{\sigma_{1}i_{1}}\varphi_{\sigma_{2}i_{2}}\varphi_{\sigma_{3}i_{3}}\ldots\varphi_{\sigma_{k}i_{k}}\left|F\right\rangle,\qquad i_{l}>i_{l-1}+1$ (29)
is an exact ground state. But these non-overlapping magnon states do not exhaust all possible ground states. For example, let us consider the overlapping magnon state
$\Omega_{\sigma_{1}}(i)\varphi_{\sigma_{2}i}\left|F\right\rangle$ (30)
where
$\Omega_{1}(i)$ $=$ $\sigma_{i-1}^{-}+S^{-}(i-1,i)+\tau_{i+1}^{-}$ (31)
$\Omega_{2}(i)$ $=$ $S^{-}(i-1,i)[\sigma_{i-1}^{-}+\frac{1}{2}S^{-}(i-1,i)+\tau_{i+1}^{-}]$ (32)
$\Omega_{3}(i)$ $=$ $S^{2-}(i-1,i)[\sigma_{i-1}^{-}+\frac{1}{3}S^{-}(i-1,i)+\tau_{i+1}^{-}]$ (33)
and $S^{-}(i-1,i)$ is the total lowering operator of ($i-1)$-th and $i$-th diamonds. It can be checked that the functions of the type (30) are the eigenstates with zero energy of the local Hamiltonians $\hat{H}_{i-1}$, $\hat{H}_{i}$, and of all the others.
We conjecture that in similarity to the OBC case the problem of counting the number of ground states is reduced to the one for a system of $n$ non-interacting spins $\frac{3}{2}$. Then, the number of ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is determined by equations
$k$ $=$ $n_{1}+2n_{2}+3n_{3}$
$l$ $=$ $n_{1}+n_{2}+n_{3}$ (34)
where $(n-l),n_{1},n_{2},n_{3}$ are numbers of spins $s=\frac{3}{2}$ with $s_{z}=\frac{3}{2}$, $\frac{1}{2}$, $-\frac{1}{2}$, $-\frac{3}{2}$, correspondingly. The solution of a combinatorial problem (34) gives the number of ground states $W(n,k)$ in form
$W(n,k)=\sum_{l}\sum_{m}C_{n}^{l}C_{l}^{k-l-m}C_{k-l-m}^{m}$ (35)
Using (35) we can count up the total ground state degeneracy $W(n)$
$W(n)=\sum_{k}W(n,k)=4^{n}$ (36)
Results (35) and (36) have been confirmed by ED calculation of finite diamond chains. The residual entropy per spin in accordance with Eq.(18) is
$s_{0}=\frac{2}{3}\ln 2$ (37)
Figure 8: Dependence of the energy gap on the system size for distorted diamond case shown in Fig.3b, calculated numerically for $N=12,15,18,24$.
The excitation spectrum of the distorted diamond chain at the transition point is gapped. The gap in the one-magnon spectrum $\Delta E_{1}$, given by Eq.(9) for the diamond in Fig.3b, is $\Delta E_{1}=1/3$. Numerical calculations show that the minimal gap is lower, but still finite $\Delta E\simeq 0.22$ (see Fig.8). Therefore, for $T<\Delta E$ the system properties are determined by the contribution of the ground state manifold or, equivalently, it is given by the one for the system of $n$ non-interacting spins $\frac{3}{2}$.
Figure 9: Temperature dependencies of the specific heat (solid line), entropy (dashed line) and the product of magnetic susceptibility and temperature, $\chi T$, (dotted line) for the distorted diamond case shown in Fig.3b, calculated numerically for $N=18$. For better visibility the specific heat is multiplied by a factor of 3, $3C(T)$.
The thermodynamic properties of the distorted diamond chain shown in Fig.3b are demonstrated in Fig.9. As it is seen in Fig.9, the entropy is equal to its residual value (37) at low temperature $T<\Delta E$ and tends to its high temperature limiting value $\ln 2$. The specific heat $C(T)$ has one maximum with a wide left shoulder. The susceptibility per spin for $T<\Delta E$ behaves as for the system of non-interacting spins $\frac{3}{2}$: $\chi=\frac{5}{12T}$. For high temperature the susceptibility tends to $\chi=\frac{1}{4T}$.
Figure 10: Magnetization curve as a function of $h/T$ for distorted diamond case shown in Fig.3b, calculated numerically for $N=18$ and $T=0.1$ (dashed line) and $T=1$ (dotted line). Magnetization curve (38) is shown by solid line.
The magnetization per spin of the system of non-interacting spins $\frac{3}{2}$ is given by
$m=\frac{3\sinh(\frac{3h}{2T})+\sinh(\frac{h}{2T})}{6\cosh(\frac{3h}{2T})+6\cosh(\frac{h}{2T})}$ (38)
and is shown in Fig.10 by solid line. As one can see in Fig.10, the magnetization curve for $T=0.1$, which is lower than the energy gap, perfectly coincides with Eq.(38). But the magnetization curve for relatively high temperature $T=1$ substantially deviates from Eq.(38).
II.3 Ideal diamond chain
For the ideal diamond chain the interactions $J_{i}$ are $J_{1}=J_{3}=J_{4}=J_{5}=-1$ and $J_{2}=1$ (see Fig.3c). In this special case there is a local conservation of the composite spin $\mathbf{L}_{i}=(\mathbf{\xi}_{i}+\mathbf{\eta}_{i})$ on the vertical diagonals of the diamonds. The composite spin is a conserved quantity with quantum spin number $L_{i}=0$ or $L_{i}=1$. Then the model reduces to the spin chain with alternating spins $s=\frac{1}{2}$ and composite spins $L$, and it is described by the Hamiltonian
$\hat{H}=-\sum[\mathbf{L}_{i}\cdot(\mathbf{s}_{i}+\mathbf{s}_{i+1})-1]+\frac{1}{2}(\mathbf{L}_{i}^{2}-2)$ (39)
where $\mathbf{L}_{i}$ and $\mathbf{s}_{i}$ are composite spin and spin $s=\frac{1}{2}$, respectively, on $i$-th diamond. The constants in Eq.(39) are chosen so that the ground state energy of Hamiltonian (39) is zero.
All ground states of (39) can be enumerated as follows. First, we choose a definite configuration of $k$ singlets $L=0$ located in diamonds $\{i_{1},i_{2},\ldots i_{k}\}$. The number of such configurations with $k$ singlets is $C_{n}^{k}$. We note that the singlet state on the diagonal of the diamond is an exact state for this diamond, independent of the spin states on the left and the right sites of it. Effectively, each singlet $L_{i}=0$ cuts the chain and creates an open boundary at this place. Therefore, for the chosen configuration of singlets $\{i_{1},i_{2},\ldots i_{k}\}$, we have $k$ sections of open chains, located between spins with $L=0$. The ground state of each of these open chains is the ferromagnetic state with all possible projections $S^{z}$, which can be written as
$\psi\left(S_{ij}^{z}\right)=\left|S_{max}(i,j),S_{ij}^{z}\right\rangle$ (40)
where the total spin of the ferromagnetic state of the open diamond chain between the singlets on diagonals of $i$-th and $j$-th diamonds is $S_{max}(i,j)=\frac{3}{2}(j-i)-1$ and the spin projection can be $S_{ij}^{z}=-S_{max},\ldots S_{max}$. Consequently, the number of multiplets of the ferromagnetic state between the singlets on $i$-th and $j$-th diamonds is $Q_{ij}=3(j-i)-1$.
Thus, the set of ground state wave functions for the configuration of singlets $\{i_{1},i_{2},\ldots i_{k}\}$ can be written as
$\Psi\left(S_{i_{1}i_{2}}^{z},S_{i_{2}i_{3}}^{z},\ldots S_{i_{k}i_{1}}^{z}\right)=\left|s_{i_{1}}\right\rangle\psi\left(S_{i_{1}i_{2}}^{z}\right)\left|s_{i_{2}}\right\rangle\psi\left(S_{i_{2}i_{3}}^{z}\right)\ldots\left|s_{i_{k}}\right\rangle\psi\left(S_{i_{k}i_{1}}^{z}\right)$ (41)
where $\left|s_{i}\right\rangle$ means the singlet state on diagonal of $i$-th diamond. Therefore, the set of wave functions (41) contains $Q_{i_{1}i_{2}}\cdot Q_{i_{2}i_{3}}\cdot\ldots\cdot Q_{i_{k}i_{1}}$ states. To calculate the total number of ground states one should sum up all such products for all possible singlet configurations $\{i_{1},i_{2},\ldots i_{k}\}$ with different $k$. The solution of this combinatorial problem is presented below.
First, we consider the open chain with $n$ diamonds. Let us name the total number of ground states of $n$ diamonds as $w_{n}$. Then, $w_{n}$ can be expressed as a sum
$w_{n}=\sum_{k=1}^{n}\left(3k-1\right)w_{n-k}+\left(3n+2\right)$ (42)
In this expression the first singlet is located in $k$-th diamond, therefore the remaining part contains $(n-k)$ diamonds with the total number of states $w_{n-k}$ (see Fig.11). The factor $(3k-1)$ represents the number of multiplets of the ferromagnetic state of $(k-1)$ diamonds, containing $(k-1)$ spins-$1$ and $k$ spins-$\frac{1}{2}$. The term with $k=n$ in the sum corresponds to the special case of just one singlet located on the last diamond. For this case there is just one spin-$\frac{1}{2}$ on the right side of the singlet and, therefore, we put $w_{0}=2$ for this case.
Figure 11: Scheme used for the calculation of the ground state degeneracy for an ideal diamond model.
The last term $(3n+2)$ in Eq.(42) corresponds to the case of no $L=0$ spins in the chain, and the degeneracy of this case is the number of multiplets of the ferromagnetic state of the total open chain.
The solution of Eq.(42) is
$w_{n}=9\cdot 4^{n-1}$ (43)
For the calculation of the total number of degenerate states $W_{n}$ for the cyclic chain, containing $n$ diamonds we write the equation
$W_{n}=w_{n-1}+(3n+1)+(n-1)(3n-1)+\sum_{m=1}^{n-2}m\left(3m+2\right)w_{n-m-2}$ (44)
The first term in Eq.(44) is the number of states when the last diamond (that loops the chain) contains the diagonal singlet. The second term $(3n+1)$ in Eq.(44) corresponds to the case without singlets in the chain, and the degeneracy in this case is the number of multiplets of the ferromagnetic state of the total system. The term $(n-1)(3n-1)$ corresponds to the case with one singlet on the whole chain. The sum contains terms when the last diamond belongs to the ferromagnetic section of length $m$.
The sum in (44) is calculated and gives for $W_{n}$ the exact result
$W_{n}=4^{n}+3n-1$ (45)
A similar approach can be used for the calculation of the ground state degeneracy in sectors with fixed total $S^{z}$. Let the total number of ground states of $n$ diamonds in the sector with total $S^{z}=S_{\max}^{z}-m$ be $w_{n,m}$. Then $w_{n,m}$ can be expressed as a sum
$w_{n,m}=\sum_{k=1}^{n}\sum_{l=0}^{3k-2}w_{n-k,m-l-1}+1$ (46)
Here, we follow the method of the calculation of the ground state degeneracy used in Eq.(42) and shown in Fig.11. The solutions of Eq.(46) are
$w_{n,0}$ $=$ 11 11 (47)
$w_{n,1}$ $=$ $n+1$ (48)
$w_{n,2}$ $=$ $\frac{\left(n+1\right)\left(n+2\right)}{2}$ (49)
Generally,
$w_{n,k}=W_{n,k}+2W_{n,k-1}+3W_{n,k-2}+2W_{n,k-3}+W_{n,k-4}$ (50)
where $W_{n,k}$ is given in Eq.(35) and coincides with the number of states of $n$ spins $\frac{3}{2}$ in the spin sector with total $S^{z}=S_{\max}^{z}-m$. Eq.(50) means that effectively the degeneracy of open chain of $n$ diamonds is equal to that of $(n-1)$ non-interacting spins $\frac{3}{2}$ and two ‘edge’ spins $1$.
The total number of ground states $W_{n,k}^{PBC}$ of the diamond chain with PBC in the spin sector with total $S^{z}=S_{\max}^{z}-k$ is
$W_{n,k}^{PBC}=W_{n,k}+1\qquad 0<k\leq\frac{3n}{2}$ (51)
All these results are confirmed numerically for finite ideal diamond chains.
Let us consider the point concerning the gap in the excitation spectrum for the ideal diamond chain. The energy of one-magnon excitations given by Eq.(9) for this diamond chain with PBC is
$\Delta E_{1}(q)=\frac{3\pm\sqrt{9-8\sin^{2}\frac{q}{2}}}{2}$ (52)
where the quasi-momentum is $q=2\pi l/n$. The lowest excited energy corresponds to $q=2\pi/n$ and is equal to $\Delta E_{1}=\frac{2\pi^{2}}{3n^{2}}$ at $n\gg 1$.
Now let us consider two-magnon excitations. At first, we consider the model (39) with only one spin $\mathbf{L}=0$ and all others $\mathbf{L}=1$. As mentioned earlier, the spin $\mathbf{L}=0$ cuts the chain so that the initially cyclic chain effectively transforms to the open chain with $n$ spins $\frac{1}{2}$ (including ends) and $(n-1)$ spins $1$. Let us consider one-magnon excitations in this open chain, which correspond to the two-magnon excitations of initial cyclic chain. The one-magnon spectrum of the open chain is given by Eq.(52), where $q$ is determined from the equation
$(1-\sqrt{9-8\sin^{2}q})\sin(qn)=-2\sin(q(n-2))$ (53)
The value of $q$ that corresponds to the lowest excitation, $\Delta E_{2}$, is $q=\frac{\pi}{2(n-1)}$ at $n\gg 1$. Then $\Delta E_{2}=\frac{\pi^{2}}{6n^{2}}$, which is 4x lower than $\Delta E_{1}$. Numerical calculations show that $\Delta E_{2}$ is the lowest level in the entire excitation spectrum. Thus, the spectrum of the ideal diamond chain is gapless in the transition point and behaves as $\Delta E\sim n^{-2}$.
As well as for the distorted diamond chain at the transition point, magnetization in the ideal diamond model is zero at $T=0$ and it undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=0$ at the transition point. The susceptibility and the specific heat of the ideal diamond model has been calculated in [27] for a chain with 120 spins. It was shown that the susceptibility diverges as $\chi\sim\frac{1}{T}$ and $\chi T=\frac{5}{12}$ at $T\to 0$ as for the distorted diamond chain. The specific heat $C(T)$ calculated in [27] has one maximum.
Deviation from the transition point leads to the ferromagnetic ground state for $J_{2}<1$ and for $J_{2}>1$ the ground state is $2^{n}$ degenerate and includes the states with $S_{tot}$ for $0\leq S_{tot}\leq\frac{n}{2}$. In terms of the Hamiltonian (39) the ground state corresponds to all $L_{i}=0$ and $n$ free spins $s_{i}$. There is a finite gap to the state with $S_{tot}=\frac{n}{2}+1$, caused by the destruction of one of the singlets $L_{i}=0$.
II.4 Diagonal state case
Let us consider the diamond model for which the the condition (10) is satisfied but (14) is not (a particular case is shown in Fig.3d). As it was noted above, the exact one-magnon states are localized on vertical diagonals of the diamond chain and they are given by Eq.(11). The exact ground states in the spin sector $S_{tot}=S_{\max}-k$ $\ (0\leq k\leq n)$ can be written as a product of $k$ functions
$\hat{\phi}_{i_{1}}\hat{\phi}_{i_{2}}\ldots\hat{\phi}_{i_{k}}\left|F\right\rangle$ (54)
where the localized states $\hat{\phi}_{i}$ can be located on neighboring diamonds.
The number of ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is equal to
$W(n,k)$ $=$ $\sum_{i=0}^{k}C_{n}^{i}\qquad 0\leq k<n$ (55)
$W(n,k)$ $=$ $2^{n}\qquad n\leq k\leq\frac{3n}{2}$ (56)
The total number of ground states $W(n)$ is
$W(n)=2^{n}(2n+1)$ (57)
The ground state degeneracy in this type of the diamond chain is exponentially large as well. In accordance with Eq.(18), this leads to the residual entropy $s_{0}=\frac{1}{3}\ln 2$ at $n\gg 1$.
Figure 12: Temperature dependencies of the specific heat (solid line), entropy (dashed line) and the product of magnetic susceptibility and temperature, $\chi T$, (dotted line) for the diagonal state case shown in Fig.3d, calculated numerically for $N=18$. For better visibility the specific heat is multiplied by a factor of 3, $3C(T)$.
The thermodynamic properties of the diagonal state chain shown in Fig.3d are demonstrated in Fig.12. As it is seen in Fig.12, the entropy tends to finite value defined by Eqs.(18),(57) at low temperatures $T<0.03$, and approaches to $\ln 2$ at high temperatures. The specific heat $C(T)$ has two maxima. The magnetic properties of the diagonal state chain is similar to the general case. The susceptibility diverges at $T\to 0$ faster than $1/T$. The magnetization curve in the thermodynamic limit has a form given by Eq.(19). Therefore, the diamond chain of this type is magnetically ordered at the transition point at $T=0$ and the magnetization in similarity to the general case of the diamond chain undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point.
Figure 13: Dependence of the energy gap on the system size for diagonal state case shown in Fig.3d, calculated numerically for $N=12,18,24,30$ (exact diagonalization and Lanczos algorithm). Dotted line is a fitting curve $\Delta E=\frac{3.12}{n^{2}}-\frac{3.8}{n^{3}}$.
The gap in one-magnon excitation spectrum (9) is equal to $\Delta E_{1}=\frac{5\pi^{2}}{6n^{2}}$ for the chain consisting of diamonds shown in Fig.3d. This means that the excitation gap goes down as $n^{-2}$ or faster at $n\gg 1$. The numerical calculations of the gap shown in Fig.13 confirm this law, and give $\Delta E=\frac{3.12}{n^{2}}$.
III Summary
We have studied the ground state of the spin-$1/2$ Heisenberg diamond chain with competing F and AF exchange interactions at the transition point between the ferromagnetic and other (singlet or ferrimagnetic) ground state phases. The ground state properties of this model at the transition point are highly nontrivial. The ground state consists of the localized magnons and special magnon complexes which form macroscopically degenerate ground state. We consider four types of diamond chains. One of them is the diamond chain with only one-magnon localized states. The properties of this model are similar to those of the F-AF delta-chain. In particular, the ground state degeneracy is $W(n)\simeq 2^{n}$ and the residual entropy is $\frac{1}{3}\ln 2$. The diamond chain of this type is magnetically ordered at $T=0$ and the magnetization undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point.
The most striking feature of the second considered model, the distorted diamond chain, is the existence of two- and three-magnon localized states along with the conventional one-magnon localized states. These states form flat bands in two- and three magnon spectrum and they are ground states as well. It turns out that the number of ground states in the sectors with fixed total spin projection $S^{z}$ coincides with the one for the system of $n$ non-interacting spins $\frac{3}{2}$. We have checked this fact by numerical calculations of finite chains. The localized magnon states in the distorted diamond chain lead to the ground state degeneracy $W(n)=4^{n}$, which exceeds exponentially the one in the model with only one-magnon localized states. As a result, the residual entropy per spin is $s_{0}=\frac{2}{3}\ln 2$. The ideal diamond chain has the same residual entropy. For the ideal diamond chain, analytical results have been obtained for the number of the ground states in different sectors of $S^{z}$ and they are confirmed by numerical calculations of finite chains. The spectrum of excitations in the three types of diamond chains is gapless and it is gapped in the distorted diamond chain. The magnetization is zero at $T=0$ at the transition point for the distorted and ideal diamond chains and, therefore, there is a magnetization jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=0$ at the transition point. The susceptibility per spin $\chi$ diverges as $\chi\sim 1/T$ and the product $\chi T$ is $\chi T=\frac{5}{12}$ at $T=0$. The main features of four cases of the diamond model are summarized in the Table.1.
Table 1: Properties of four cases of diamond model.
General Distorted Ideal Diagonal
Ground state degeneracy $2^{n}+(n+4)C_{n}^{n/2}$ $4^{n}$ $14^{n}+3n-$ 1 $(2n+1)2^{n}$
Residual entropy per spin $\frac{1}{3}\ln 2$ $\frac{2}{3}\ln 2$ $\frac{2}{3}\ln 2$ $\frac{1}{3}\ln 2$
Lowest excitation, $\Delta E$ gapless, $\sim e^{-cn}$ gapped gapless, $\sim n^{-2}$ gapless, $\sim n^{-2}$
Low-T susceptibility, $\chi T$ diverges $\frac{5}{12}$ $\frac{5}{12}$ diverges
Specific heat two maxima one maximum one maximum two maxima
In this paper we consider the isotropic diamond chain on the boundary of the ferromagnetic ground state. This model can be extended to the case of anisotropic interactions similar to those developed for F-AF delta chain in Ref.[32, *anis2, *anis3] (details will be presented elsewhere).
It is known that spin systems with the macroscopic degenerate ground state show enhanced magnetocaloric effect. From this point of view the considered diamond chain can serve as a beacon for the search for materials for low-temperature cooling. Of course, macroscopic degeneracy occurs for certain relations between exchange interactions, which can hardly be satisfied exactly in real compounds. However, some important features of the model remain if the values of interactions are close to the critical ones. Recent progress in the synthesis of diamond chain compounds allows us to expect that compounds with parameters close to those given in this work can be synthesized in the near future.
Acknowledgements.
The numerical calculations were carried out with use of the ALPS libraries [35]. This work was financially supported by the Ministry of Sciences and Higher Education, Russian Federation (Research theme state registration number 122041400110-4).